OpenStudy (anonymous):

What is the percent of decrease from 320 to 32?

OpenStudy (anonymous):

is it 90%

OpenStudy (texaschic101):

percent decrease = (older - newer) / older, then multiply that by 100. percent decrease = (320 - 32) / 320 percent decrease = 288/320 percent decrease = 0.9 0.9 x 100 = 90% you are correct